OpenStudy (amistre64):

Why is 0! = 1 by definition; but 0^0 is an indeterminate form?

OpenStudy (anonymous):

! <------Because of this...maybe

OpenStudy (amistre64):

im curious because x^0 is bad form since it can result in 0^0; but why not define it as 1 with the rest of normailty :)

OpenStudy (anonymous):

even I am curious about that

OpenStudy (anonymous):

don't change it to 1 ! it will mess up a lot of math ppl already did in the past. just keep the status quo!

OpenStudy (anonymous):

but imran told me 1^0 is 1 ...2^0 is 1 and 2^1 is 2 so 0^0 is indeterminate

OpenStudy (amistre64):

yep; 0^0 is a bad an apple as 1/0 ....

OpenStudy (amistre64):

but if you get 0!, well, thats just fine lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Actually, 0! = 1 by contemporary tradition. Many people claim that it is 0 via proof. A lot of well known math and applied physics professors have claimed that 0! = 0 and keep 0! = 1 simply because our formulas break down without it true.

OpenStudy (anonymous):

Theres an awesome site that has professor proofs of 0! = 0 let me see if i can find it

OpenStudy (amistre64):

lol ... google 0^0 = 1, i love it

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

and we had doubts today about wolfram Ishaan :)

OpenStudy (amistre64):

what did wolfram do?

OpenStudy (anonymous):

he was smart ! that tells us what we were.

OpenStudy (anonymous):

Indeterminate http://www.wolframalpha.com/input/?i=0^0

OpenStudy (anonymous):

Lol

OpenStudy (amistre64):

i seen that as well

OpenStudy (amistre64):

it began me thinking with why we have 2 rules for derivatives; d\dx (x) = 1 d/dx (x^n) = n*x^(n-1)

OpenStudy (amistre64):

d/dx (x^1) = 1*x^1-1 = 1*x^0 ..right?

OpenStudy (amistre64):

but this goes bad when x=0 .... for some reason that cant be defined

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

http://www.adonald.btinternet.co.uk/Factor/Zero.html

OpenStudy (anonymous):

interesting i guess

OpenStudy (amistre64):

its a cute story :)

OpenStudy (amistre64):

but check this out :) http://www.wolframalpha.com/input/?i=y+%3D+x^0+at+x%3D0

OpenStudy (anonymous):

That is so confusing

OpenStudy (anonymous):

That's why it's called indeterminant - nobody know what the heck to do with it.

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

But seriously I am very confused now...

OpenStudy (anonymous):

hold on. there is lots of controversy (ok a little) still about what \[0^0\] is of course as a statement about limits it is indeterminate, but as a symbol many define it to be 1. those are different issues

OpenStudy (amistre64):

wolfram seems to be schismed about it as well

OpenStudy (amistre64):

0^0= bad form http://www.wolframalpha.com/input/?i=0^0 but, f(x)=x^0 ; f(0)=1 http://www.wolframalpha.com/input/?i=y+%3D+x^0+at+x%3D0

OpenStudy (anonymous):

some sources here http://mathforum.org/dr.math/faq/faq.0.to.0.power.html

OpenStudy (amistre64):

we seem to be in my own little universe ... all i see is my questions, and i cry ...

OpenStudy (anonymous):

There once was a student from Trinity, Tried to take the square root of infinity. Whilst counting the digits Was seized with the fidgets. Gave up math and took up divinity.

OpenStudy (anonymous):

only one of two clean ones i know

OpenStudy (amistre64):

limericks eh ... that was a good read.

OpenStudy (amistre64):

i like the 0^x is rather unimportant; yet if 0^x = 0 for the most part; then why would we attribute it to 0^0 = 1 instead of 0 if x^0 is 1 for the most part, how can we define it as 0 as well. and of course they have different limits and such

OpenStudy (anonymous):

not particularly one of the best but one of the few i would post here